OpenStudy (anonymous):

The sum of two nonegative numbers is 20. Find the numbers if...

OpenStudy (anonymous):

a) the sum of their squares is a large as possible; as small as possible? b) one number plus the square root of the other is a large as possible; as small as possible?

OpenStudy (jennychan12):

i only know how to do this with optimization (calculus) x^2+y^2 =M x+y = 20 Find y in terms of x and plug it in, then take the derivative.

OpenStudy (anonymous):

I got y = 10

OpenStudy (anonymous):

so that's the minimum? but i don't know how to find the largest possible answer

OpenStudy (anonymous):

no, i did it your previous way lol. I put the primary equation as x^2+y^2= S and the secondary equation as x+y=20 but idk if that's correct.

OpenStudy (jennychan12):

yeah, that's correct, then solve for one of the variables and plug it into the S equation

OpenStudy (anonymous):

the S equation or the S' equation?

OpenStudy (hunus):

the S equation

OpenStudy (jennychan12):

plug it into the S equation, simplify then take the derivative (S')

OpenStudy (anonymous):

Oh i figured it out lol!!! But i don't know how to solve for part b :(

OpenStudy (jennychan12):

x+sqrt(y) = M try that?

OpenStudy (jennychan12):

using x+y = 20 plug in, same things...? dunno if it'll work, but....